OpenStudy (anonymous):

What are the next three terms of the sequence 1, -3, 9, -27,…? A. -54, 162, -486 B. -81, 243, -729 C. 81, -243, 729 D. 57, -162, 486

OpenStudy (tkhunny):

Ridiculous problem. Finite sequences have infinitely many "next three terms". ANYTHING you can justify is an acceptable answer. 1, -3, 9, -27, 0, -2, 5 -- That works. 1, -3, 9, -27, 28, -29, 5 -- That works. 1, -3, 9, -27, \(\sqrt{5}\), \(\pi\), e -- That works. What say you? Give us a clue what YOU might think are three acceptable values and justify your response.

OpenStudy (anonymous):

Well, I thought it was 57, -162, 486, but I really wasn't sure, so I came on here for some help..

OpenStudy (tkhunny):

It's not a matter of being sure. It's a matter of justifying your answer. How did you get those three values?

OpenStudy (anonymous):

If you're having trouble with this, try breaking it down. The sequence itself has 4 values in it already. What did the 1 do to become -3?

OpenStudy (tkhunny):

Generally, it's a good idea to look for common differences or common ratios. Subtract each successive pair. What do you get? Divide each successive pair. What do you get?

OpenStudy (anonymous):

i really don't understand

OpenStudy (loser66):

1*(-3) =?

OpenStudy (anonymous):

-3

OpenStudy (loser66):

Now, second term -3*(-3) =?

OpenStudy (anonymous):

9

OpenStudy (loser66):

Now, third term 9*(-3) =?

OpenStudy (anonymous):

-27

OpenStudy (loser66):

can you see your answers are the term of the given sequence??

OpenStudy (anonymous):

oh, yeah! Thank you so much! I get it now :)

OpenStudy (loser66):

so, next will be???

OpenStudy (anonymous):

81

OpenStudy (anonymous):

then it'll be -243

OpenStudy (loser66):

Yes!! look at your options, which is correct.

OpenStudy (anonymous):

c! Thanks again :)

OpenStudy (loser66):

No need to do next, since there is only 1 option which starts at 81, It is C

OpenStudy (loser66):

YOu got it. hehehe...